# A Fourier-RKHS approach for characterizing equivalent Gaussian distributions for stationary processes on homogeneous spaces

### v1

@@ 1.5.1 Positive-definite functions
Let $G$ be a locally compact group. In general, we write $g\cdot\tilde{g}=g\tilde{g}$ for the group operation and $e=gg^{-1}$ identifies the identity element on $G$. A continuous function $\varphi\colon G\to\mathbb{C}$ is said to be positive-definite if for any $n\in\mathbb{N}$ and $g_{1},\dotsc,g_{n}\in G$,
@@ 1.5.2 Positive-definite functions on LCA groups
$\chi(g+\tilde{g})=\chi(g)\chi(\tilde{g}),\quad g,\tilde{g}\in G.$
In particular, $\chi(0)=1$ and $\chi(-g)=\overline{\chi(g)}$, $g\in G$. The following is known as Bochner’s theorem [Rudin]. A continuous function $\phi\colon G\to\mathbb{C}$ is positive-definite if and only if,
$\phi(g)=\int_{G^{*}}\chi(g)\mu(d\chi),\quad g\in G,$ (13)
@@
If $G=\mathbb{R}^{d}$, the dual group $G^{*}$ is isomorphic to $\mathbb{R}^{d}$ with isomorphism $\lambda\mapsto\operatorname{exp}({i\mkern 1.0mu}\langle\cdot,\lambda\rangle)$. In particular, any continuous character $\chi$ of $\mathbb{R}^{d}$ is given by $\chi(\cdot)=\chi_{\lambda}(\cdot)=\operatorname{exp}({i\mkern 1.0mu}\langle\cdot,\lambda\rangle)$ for some $\lambda\in\mathbb{R}^{d}$. Upon the identification $T\colon G^{*}\to\mathbb{R}^{d}$, $T(\chi_{\lambda})=\lambda$, it follows from (13) that any continuous positive-definite function $\phi\colon\mathbb{R}^{d}\to\mathbb{C}$ admits a representation $\phi(x)=\int_{\mathbb{R}^{d}}\!\operatorname{exp}({i\mkern 1.0mu}\langle x,\lambda\rangle)F(d\lambda)$ where $F$ is the pushforward of $\mu$ by $T$.
@@ 1.5.3 Equivalent Gaussian distributions on LCA groups
Then, the following characterization of equivalent Gaussian distributions on LCA groups is due to Grenander [Grenander] (cf. [Chatterji_Mandrekar]). For $P_{1}$ and $P_{2}$ to be equivalent on $\sigma_{T}(X)$ it is necessary and sufficient that the nonatomic parts of $\mu_{1}$ and $\mu_{2}$ agree and $\mu_{1}$ and $\mu_{2}$ share the same set of atoms $\{a_{n}\colon n\in\mathbb{N}\}$ s.t.
@@ 1.6.1 Spherical characterization
The latter representation is due to Godement [Godement1958] (cf. the earlier work of Gelfand [Gelfand]).
@@ 1.6.2 Translation to invariant kernels on homogeneous spaces
Given the choices for $G$ and $K$, besides the requirements for $R_{\ell}$ (resp. $\varphi_{\ell}$), the only assumption underlying (15) is that $(G,K)$ is a Gelfand pair. Specifically, if $G$ is a Lie group, any (Riemannian) symmetric pair $(G,K)$ is a Gelfand pair. This result is due to Gelfand [Gelfand] (cf. [Helgason], p. 408). In the context of invariant stochastic processes, having in mind the appearance of (15), further reading is given in the book by Malyarenko [Malyarenko] (cf. Yaglom [YaglomInvariant]).
@@
If $G=SO(d)$, $d\in\mathbb{N}$, the covariance function given in Example 1.5, $R_{\ell}(s,t)=\psi_{\ell}(\Delta(s,t))$, is $G$ invariant — its representation based on spherical harmonics is an instance of (15) with Gelfand pair (SO(d),SO(d1))(SO(d),SO(d-1)).
@@ 1.6.3 Equivalent Gaussian distributions on homogeneous spaces
Given $\zeta\in\mathscr{L}$, it is always possible to find a corresponding (spherical) irreducible unitary representation $\pi_{\zeta}$ on a Hilbert space $H_{\zeta}$ with cyclic vector $u_{\zeta}$. That is, for any $g\in G$, $\zeta(g)=\langle\pi_{\zeta}(g)u_{\zeta},u_{\zeta}\rangle_{\zeta}$, where $\langle\cdot,\cdot\rangle_{\zeta}$ denotes the inner product on $H_{\zeta}$ (see for instance Theorem 3.4, Chapter IV, §3, in [Helgason]). If $H_{\zeta}$ is finite dimensional, $d(\zeta)$ denotes the dimension of $H_{\zeta}$, otherwise $d(\zeta)=\infty$. The following is due to Chow [chow1, chow2] [^1] Chow’s result is stated more generally. It allows for less symmetric $T$, such as $G$-spaces possessing a dense orbit. We have formulated the statement in the context of Gelfand pairs $(G,K)$.. Under the assumption of (15), $P_{1}\equiv P_{2}$ on $\sigma_{T}(X)$ if and only if
@@
The proof of Chow is based on Feldman’s [Feldman1958] characterization of equivalent Gaussian distributions. In the following, we will give an RKHS interpretation of Chow’s result.
@@ 2 Equivalent Gaussian distributions on homogeneous spaces: A Fourier-RKHS approach
Let $T=G/K$, where $G$ and $K$ are as in the previous section. In particular, $G$ is unimodular locally compact and $K$ a compact subgroup of $G$ s.t. $(G,K)$ is a Gelfand pair. Assume that $R_{\ell}$ is $G$-invariant with $K$-invariant version $\varphi_{\ell}$ that is continuous at the identity. Let $\varphi_{\ell}\in L^{1}(G)\cap L^{2}(G)$ — or equivalently, $R_{\ell}(\cdot,t)\in L^{1}(T)\cap L^{2}(T)$ for any $t\in T$. Given $f\in L^{1}(T)$, we write $\hat{f}(\zeta)$ for the Fourier transform of $f$ at $\zeta\in\mathscr{L}$ (cf. [Wolf], p. 196). Let $\mathcal{H}(\ell)$ be composed of continuous functions $f\in L^{1}(T)\cap L^{2}(T)$ which are of the form
@@
where $\mu_{\scriptscriptstyle P}$ is the Plancherel measure for $(G,K)$ on $\mathscr{L}$. Since $R_{\ell}$ is real-valued, we accept the notational convenience $\operatorname{Re}f=f$. Then, if we equip $\mathcal{H}(\ell)$ with inner product
@@
it follows from the Plancherel theorem (cf. Corollary 9.6.6 in [Wolf]) that $\mathcal{H}(\ell)$ is a Hilbert space. In particular, for any $t\in T$, $R_{\ell}(\cdot,t)\in\mathcal{H}(\ell)$ and $f(t)=\langle f,R_{\ell}(\cdot,t)\rangle_{\mathcal{H}(\ell)}$, $f\in\mathcal{H}(\ell)$. This is summarized in the following lemma.
@@
For $\varphi_{\ell}\in L^{1}(G)\cap L^{2}(G)$ and $G$ not compact, $P_{1}\equiv P_{2}$ on $\sigma_{T}(X)$ if and only if $\mu_{1}=\mu_{2}$. To see it, notice that for any $v,w\in G$, the difference $R_{2}-R_{1}](vK,wK)$ is given by $\varphi_{2}-\varphi_{1}](v^{-1}w)$. We know that $P_{1}\equiv P_{2}$ on $\sigma_{T}(X)$ is equivalent to $R_{2}-R_{1}\in H_{T\times T}(R_{1}^{2\otimes})$. By the lemma, the latter is equivalent to $R_{2}-R_{1}\in L^{2}(T\times T)$, which is true if and only if $\varphi_{1}=\varphi_{2}$. Or equivalently, $\mu_{1}=\mu_{2}$ (by uniqueness of the measures $\mu_{1}$ and $\mu_{2}$).
@@
We recall that if $G$ is compact, $d(\zeta)<\infty$, $\zeta\in\mathscr{L}$, and the Plancherel measure $\mu_{\scriptscriptstyle P}$ is purely atomic (cf. Chapter 5 in [Wolf]). Also, by Mercer’s theorem (see for instance [RKHSBook]), $\varphi_{\ell}\in L^{2}(G)$. In particular, we are in the setting of the lemma. By construction of the Plancherel measure, since $\varphi_{\ell}$ is bi-invariant, positive-definite and a member of $C(G)\cap L^{1}(G)$, we have that $\mu_{\ell}(d\zeta)=\hat{\varphi}_{\ell}(\zeta)\mu_{\scriptscriptstyle P}(d\zeta)$. In the latter, $\hat{\varphi}_{\ell}(\zeta)$ is the spherical transform of $\varphi_{\ell}$ at $\zeta$ according to Theorem 9.4.1 in [Wolf]. Therefore, $\mu_{\ell}$ is purely atomic. Also, by uniqueness of the Plancherel measure on $\mathscr{L}$, $\mu_{1}$ and $\mu_{2}$ have the same set of atoms $\{a_{n}\colon n\in\mathbb{N}\}$. In particular, up to $\mu_{\scriptscriptstyle P}$ measure zero,
@@
By the lemma, $P_{1}$ and $P_{2}$ are equivalent on $\sigma_{T}(X)$ if and only if
@@
That is, $P_{1}\equiv P_{2}$ on $\sigma_{T}(X)$ if and only if Chow’s (ii) is satisfied.
@@ 3 On ML-inference for centered Gaussian processes
relates to the feasibility to estimate $\theta_{0}$ consistently. In particular, a violation of (20) can lead to inconsistent ML estimators [zhang, zhang2005towards, anderes2010consistent, bevilacqua2019estimation, Bachoc2021]. Building on a classical likelihood argument [wald], it is possible to obtain strongly consistent ML covariance parameter estimators for families of orthogonal Gaussian distributions (cf. [zhang, hediger]). The argument relies on the fact that under the separability condition (20), the Radon-Nikodym derivative $\nicefrac{{p_{\theta}^{n}(Y_{n})}}{{p_{\theta_{0}}^{n}(Y_{n})}}$ converges to zero with $P_{\theta_{0}}$ probability one whenever $\theta\neq\theta_{0}$ (cf. Theorem 1 on p. 442 in [StochasticProcesses]). Then, under suitable conditions on the parameter space and the likelihood function, the latter convergence can be strengthened to a $P_{\theta_{0}}$ a.s. uniform convergence on $\Theta$ outside a small neighborhood of $\theta_{0}$. Regarding weak consistency, the reader is referred to [sweeting1980uniform, mardia1984maximum, bachoc2014asymptotic]. In particular, in the latter two, weakly consistent estimators are obtained under the assumption that the distance between coordinates from $T$ is uniformly bounded away from zero.

### v2

@@ 1.5.1 Positive-definite functions
Let $G$ be a locally compact group. In general, we write $g\cdot g^{\prime}=gg^{\prime}$ for the group operation and $e=gg^{-1}$ identifies the identity element on $G$. A continuous function $\varphi\colon G\to\mathbb{C}$ is said to be positive-definite if for any $n\in\mathbb{N}$ and $g_{1},\dotsc,g_{n}\in G$,
@@ 1.5.2 Positive-definite functions on LCA groups
$\chi(g+g^{\prime})=\chi(g)\chi(g^{\prime}),\quad g,g^{\prime}\in G.$
In particular, $\chi(0)=1$ and $\chi(-g)=\overline{\chi(g)}$, $g\in G$. The following is known as Bochner’s theorem [39]. A continuous function $\varphi\colon G\to\mathbb{C}$ is positive-definite if and only if,
$\varphi(g)=\int_{G^{*}}\chi(g)\mu(d\chi),\quad g\in G,$ (13)
@@
If $G=\mathbb{R}^{d}$, the dual group $G^{*}$ is isomorphic to $\mathbb{R}^{d}$ with isomorphism $\lambda\mapsto\operatorname{exp}({i\mkern 1.0mu}\langle\cdot,\lambda\rangle)$. In particular, any continuous character $\chi$ of $\mathbb{R}^{d}$ is given by $\chi(\cdot)=\chi_{\lambda}(\cdot)=\operatorname{exp}({i\mkern 1.0mu}\langle\cdot,\lambda\rangle)$ for some $\lambda\in\mathbb{R}^{d}$. Upon the identification $\Lambda\colon G^{*}\to\mathbb{R}^{d}$, $\Lambda(\chi_{\lambda})=\lambda$, it follows from (13) that any continuous positive-definite function $\varphi\colon\mathbb{R}^{d}\to\mathbb{C}$ admits a representation $\varphi(x)=\int_{\mathbb{R}^{d}}\!\operatorname{exp}({i\mkern 1.0mu}\langle x,\lambda\rangle)F(d\lambda)$ where $F$ is the pushforward of $\mu$ by $\Lambda$.
@@ 1.5.3 Equivalent Gaussian distributions on LCA groups
Then, the following characterization of equivalent Gaussian distributions on LCA groups is due to Grenander [23] (cf. [11]). For $P_{1}$ and $P_{2}$ to be equivalent on $\sigma_{T}(X)$ it is necessary and sufficient that the nonatomic parts of $\mu_{1}$ and $\mu_{2}$ agree and $\mu_{1}$ and $\mu_{2}$ share the same set of atoms $\{a_{n}\colon n\in\mathbb{N}\}$ s.t.
@@ 1.6.1 Spherical characterization
The latter representation is due to Godement [21] (cf. the earlier work of Gelfand [17]).
@@ 1.6.2 Translation to invariant kernels on homogeneous spaces
Given the choices for $G$ and $K$, besides the requirements for $R_{\ell}$ (resp. $\varphi_{\ell}$), the only assumption underlying (15) is that $(G,K)$ is a Gelfand pair. Specifically, if $G$ is a Lie group, any (Riemannian) symmetric pair $(G,K)$ is a Gelfand pair. This result is due to Gelfand [17] (cf. [26], p. 408). In the context of invariant stochastic processes, having in mind the appearance of (15), further reading is given in the book by Malyarenko [31] (cf. Yaglom [50]).
@@
If $G=SO(d)$ is the rotation group in dimension $d$, the covariance function $R_{\ell}$ from Example 1.5 is $G$ invariant — its representation based on spherical harmonics is an instance of (15) with Gelfand pair (SO(d),SO(d1))(SO(d),SO(d-1)).
@@ 1.6.3 Equivalent Gaussian distributions on homogeneous spaces
Given $\zeta\in\mathscr{L}$, it is always possible to find a corresponding (spherical) irreducible unitary representation $\pi_{\zeta}$ on a Hilbert space $H_{\zeta}$ with cyclic vector $u_{\zeta}$. That is, for any $g\in G$, $\zeta(g)=\langle\pi_{\zeta}(g)u_{\zeta},u_{\zeta}\rangle_{\zeta}$, where $\langle\cdot,\cdot\rangle_{\zeta}$ denotes the inner product on $H_{\zeta}$ (see for instance Theorem 3.4, Chapter IV, §3, in [26]). If $H_{\zeta}$ is finite dimensional, $d(\zeta)$ denotes the dimension of $H_{\zeta}$, otherwise $d(\zeta)=\infty$. The following is due to Chow [12, 13] [^1] Chow’s result is stated more generally. It allows for less symmetric $T$, such as $G$-spaces possessing a dense orbit. We have formulated the statement in the context of Gelfand pairs $(G,K)$.. Under the assumption of (15), $P_{1}\equiv P_{2}$ on $\sigma_{T}(X)$ if and only if
@@
The proof of Chow is based on Feldman’s [15] characterization of equivalent Gaussian distributions. In the following, we will give an RKHS interpretation of Chow’s result.
@@ 2 Equivalent Gaussian distributions on homogeneous spaces: A Fourier-RKHS approach
Let $T=G/K$, where $G$ and $K$ are as in the previous section. In particular, $G$ is unimodular locally compact and $K$ a compact subgroup of $G$ s.t. $(G,K)$ is a Gelfand pair. Assume that $R_{\ell}$ is $G$-invariant with $K$-invariant version $\varphi_{\ell}$ that is continuous at the identity. Let $\varphi_{\ell}\in L^{1}(G)\cap L^{2}(G)$ — or equivalently, $R_{\ell}(\cdot,t)\in L^{1}(T)\cap L^{2}(T)$ for any $t\in T$. Given $f\in L^{1}(T)$, we write $\hat{f}(\zeta)$ for the Fourier transform of $f$ at $\zeta\in\mathscr{L}$ (cf. [48], p. 196). Let $\mathcal{H}(\ell)$ be composed of continuous functions $f\in L^{1}(T)\cap L^{2}(T)$ which are of the form
@@
where $\mu_{\scriptscriptstyle P}$ is the Plancherel measure for $(G,K)$ on $\mathscr{L}$. Since $R_{\ell}$ is real-valued, we accept the notational convenience $\operatorname{Re}f=f$. Then, if we equip $\mathcal{H}(\ell)$ with the inner product
@@
it follows from the Plancherel theorem (cf. Corollary 9.6.6 in [48]) that $\mathcal{H}(\ell)$ is a Hilbert space. In particular, for any $t\in T$, $R_{\ell}(\cdot,t)\in\mathcal{H}(\ell)$ and $f(t)=\langle f,R_{\ell}(\cdot,t)\rangle_{\mathcal{H}(\ell)}$, $f\in\mathcal{H}(\ell)$. This is summarized in the following lemma.
@@
For $\varphi_{\ell}\in L^{1}(G)\cap L^{2}(G)$ and $G$ not compact, $P_{1}\equiv P_{2}$ on $\sigma_{T}(X)$ if and only if $\mu_{1}=\mu_{2}$. To see it, notice that for any $v,w\in G$, the difference $R_{2}-R_{1}](vK,wK)$ is given by $\varphi_{2}-\varphi_{1}](v^{-1}w)$. We know that for $P_{1}$ and $P_{2}$ to be equivalent on $\sigma_{T}(X)$ it is necessary that $R_{2}-R_{1}\in H_{T\times T}(R_{1}^{2\otimes})$. By the lemma, the latter is equivalent to $R_{2}-R_{1}\in L^{2}(T\times T)$, which is true if and only if $\varphi_{1}=\varphi_{2}$. Or equivalently, $\mu_{1}=\mu_{2}$ (by uniqueness of the measures $\mu_{1}$ and $\mu_{2}$).
@@
We recall that if $G$ is compact, $d(\zeta)<\infty$, $\zeta\in\mathscr{L}$, and the Plancherel measure $\mu_{\scriptscriptstyle P}$ is purely atomic (cf. Chapter 5 in [48]). Also, by Mercer’s theorem (see for instance [37]), $\varphi_{\ell}\in L^{2}(G)$. In particular, we are in the setting of the lemma. By construction of the Plancherel measure, since $\varphi_{\ell}$ is bi-invariant, positive-definite and a member of $C(G)\cap L^{1}(G)$, we have that $\mu_{\ell}(d\zeta)=\hat{\varphi}_{\ell}(\zeta)\mu_{\scriptscriptstyle P}(d\zeta)$. In the latter, $\hat{\varphi}_{\ell}(\zeta)$ is the spherical transform of $\varphi_{\ell}$ at $\zeta$ according to Theorem 9.4.1 in [48]. Therefore, $\mu_{\ell}$ is purely atomic. Also, by uniqueness of the Plancherel measure on $\mathscr{L}$, $\mu_{1}$ and $\mu_{2}$ have the same set of atoms $\{a_{n}\colon n\in\mathbb{N}\}$. In particular, up to $\mu_{\scriptscriptstyle P}$ measure zero,
@@
By the lemma, the equivalence of $P_{1}$ and $P_{2}$ on $\sigma_{T}(X)$ implies that
@@
Therefore, if the latter sum is infinite, $P_{1}$ and $P_{2}$ are orthogonal on $\sigma_{T}(X)$.
@@ 3 On ML-inference for centered Gaussian processes
relates to the feasibility to estimate $\theta_{0}$ consistently. In particular, a violation of (20) can lead to inconsistent ML estimators [52, 51, 1, 6, 5]. Building on a classical likelihood argument [45], it is possible to obtain strongly consistent ML covariance parameter estimators for families of orthogonal Gaussian distributions (cf. [52, 16]). The argument relies on the fact that under the separability condition (20), the Radon-Nikodym derivative $\nicefrac{{p_{\theta}^{n}(Y_{n})}}{{p_{\theta_{0}}^{n}(Y_{n})}}$ converges to zero with $P_{\theta_{0}}$ probability one whenever $\theta\neq\theta_{0}$ (cf. Theorem 1 on p. 442 in [18]). Then, under suitable conditions on the parameter space and the likelihood function, the latter convergence can be strengthened to a $P_{\theta_{0}}$ a.s. uniform convergence on $\Theta$ outside a small neighborhood of $\theta_{0}$. Regarding weak consistency, the reader is referred to [43, 32, 4]. In particular, in the latter two, weakly consistent estimators are obtained under the assumption that the distance between coordinates from $T$ is uniformly bounded away from zero.